OpenStudy (anonymous):

solve the system using the substitution method 3x-y=0 2x-y=-1

OpenStudy (anonymous):

2(3x-y=0), -3(2x-y=-1) 6x-2y=0 -6x+3y=3 y=3 6x-2(3)=0 6x=6 x=1 (1,3) Final answer